Both answers are incorrect.
Suppose the price level and value of the U.S. Dollar in year 1 are 1 and $1, respectively Instructions: Round your answers to 2 decimal places α. If the price level rises to 110 in year 2, what is the new value of the dollar? 1.10 3 b. If,instead, the price level fails to 0.55, what is the value of the dollar? 0.55 Suppose the price level and value of the US. Dollar in year 1 are 1 and $1 respectively Instructions: Round your answers to 2 decimal places a If the price level rises to 110 in year 2, what is the new value of the dollar? O Answer is complete but not entirely correct 90 0 b. It instead, the price level falls to 0.55, what is the value of the doltlar? O Answer is complete but not entirely correct 181 0
Answer #1

a) Ans: 0.91

Explanation:

The value of Dollar = 1 / Price level = 1 / 1.10 = 0.91

b) Ans: 1.82

Explanation:

The value of Dollar = 1 / Price level = 1 / 0.55 = 1.82